Cora wants to determine a 95% confidence interval for the true proportion of high school students in the area who attend their home basketball games. How large of a sample must she have to get a margin of error less than 0.03?
Answer #1

Solution,

Given that,

=  1 - = 0.5

margin of error = E = 0.03

At 95% confidence level

= 1 - 95%  

= 1 - 0.95 =0.05

/2 = 0.025

Z/2 = Z0.025  = 1.96

sample size = n = (Z / 2 / E )2 * * (1 - )

= (1.96 / 0.03 )2 * 0.5 * 0.5

= 1067.11

sample size = n = 1068